OpenStudy (mysticfury):

PLEASE HELP ITS MY LAST QUESTION WILL MEDAL Given quadrilateral LMNO ~ quadrilateral RSTU and LM = 8, NO = 12, and RS = 7, what is TU?

OpenStudy (deadpooi):

10.5?

OpenStudy (mysticfury):

@Aveline @Ms-Brains @thomaster @AloneS

OpenStudy (mathmale):

I always strongly insist on seeing a sketch of a situation like this one before commenting. Could you plot the given points, connect them with straight lines and share your work? Use the Draw utility, below.

OpenStudy (aveline):



OpenStudy (mysticfury):

yep but now how to i solve it @Aveline ? do you know?

OpenStudy (aveline):

8 times 7/8 = 7 12 times 7/8 = 84/8 84/8=10.5

OpenStudy (mysticfury):

okay thanks :D